A torsional oscillator of rotational inertia 1.6 kg⋅m2 and torsional constant 3.4 N⋅m/rad has total energy 4.3 J.

a) What is its maximum angular displacement? Express your answer in radians.

b) What is its maximum angular speed? Express your answer in radians per second.
Answer #1

PART (a):

Maximum displacement is given by:

Emax = 1/2 K θmax2

(4.3 J) = 1/2 x (3.4 N⋅m/rad) x (θmax2)

∴ θmax2) = 4.3 / 1.7

∴ θmax = 1.304 rad

-----------------------------------------------------

Also, Maximum energy = 1/2 Iω2

Emax = 1/2 Iω2

(4.3 J) = 1/2 x ( 1.6 kg⋅m2) x (ω2)

(4.3 J) = (0.8 kg⋅m2) x (ω2)

∴ ω2 = 4.3/0.8

∴ ω​​​​​​​ = 2.32 rad/s

-------------------------------